# Structural adaptation via directional regularity: rate accelerated estimation in multivariate functional data

### v2

@@ Abstract
We introduce directional regularity, a new definition of anisotropy for multivariate functional data. Instead of taking the conventional view which determines anisotropy as a notion of smoothness along a dimension, directional regularity additionally views anisotropy through the lens of directions. We show that faster rates of convergence can be obtained through a change-of-basis by adapting to the directional regularity of a multivariate process. An algorithm for the estimation and identification of the change-of-basis matrix is constructed, made possible due to the unique replication structure of functional data. Non-asymptotic bounds are provided for our algorithm, supplemented by numerical evidence from an extensive simulation study. We discuss two possible applications of the directional regularity approach, and advocate its consideration as a standard pre-processing step in multivariate functional data analysis.
@@ 1 Introduction
Anisotropy is a key concept in probability and statistics. In the context of non-parametric statistics, it roughly states that the regularity is different between dimensions. Let $(X_{m},Y_{m})$ be independent and identically distributed (iid) pairs observed under the model
$Y_{m}=f(X_{m})+\epsilon_{m},\qquad m=1,\dots,M_{0},$
with $f:[0,1]^{d}\rightarrow\mathbb{R}$, and the design points $X_{m}$’s are uniformly distributed on the hypercube $[0,1]^{d}$. The errors $\epsilon_{m}$ are assumed to be uncorrelated, centered random variables. When $f$ belongs to an anisotropic Hölder class, it is well known that under suitable assumptions on the noise, the minimax rate of estimation is $M^{-\beta/(2\beta+1)}$, where $\beta$ is given by
$\frac{1}{\beta}=\sum_{i=1}^{d}\frac{1}{\beta_{i}},$ (1)
with $\beta_{i}$ denoting the regularity along dimension $i$. The parameter $\beta$ is also known as the “effective smoothness" of $f$, since it characterizes the regularity of $f$ across different dimensions. Since $\beta^{-1}\leq\min_{i=1,\dots,d}\beta_{i}^{-1}$, anisotropy brings along better rates of convergence in the non-parametric regression setup of (1).
The regularities $\beta_{i},i=1,\dots,d$ are usually considered along the canonical basis, the domain where the data is commonly observed on. However, this can lead to misleading conclusions since anisotropy is not invariant to directions. A relatively recent body of work in non-parametric statistics aims to give consideration to the directional nature of anisotropy, called structural adaptation (Lepski, (2015)). See Samarov and Tsybakov, (2004), Lepski and Rebelles, (2020), Ammous et al., (2024) for some examples in the density model. In spatial data, this has been considered under the umbrella of fractal analysis, for example in the seminal work of Davies and Hall, (1999).
@@
Surprisingly, to the best of our knowledge, the directional nature of anisotropy has yet to be examined in multivariate functional data analysis. We aim to fill this gap by giving a rigorous definition of anisotropy through the directional regularity of a process. Through several examples, we demonstrate that performing a change-of-basis can increase the effective smoothness of a process along the canonical basis. This can lead to faster rates of convergence for ubiquitous tasks such as smoothing in functional data analysis. This change-of-basis matrix is intimately related to the notion of directional regularity that we introduce in this paper, and an algorithm is constructed to estimate it.
Functional data analysis (FDA) is a burgeoning field that deals with modern, complex data, such as those collected by sensors. For an introduction, see the textbooks Ramsay and Silverman, (2005), Horváth and Kokoszka, (2012), Kokoszka and Reimherr, (2017). FDA deals with data sets where the datum are functions, regarded as realizations of a stochastic process or random field, defined over a continuous domain. A unique feature of functional data is replication, where several sample paths are observed.
@@
The remainder of this paper is organized as follows. In Section 2 we give a formal definition of directional regularity, which is elucidated through Lemma 1 and several examples. Lemma 1, which states that the regularity of a surface is the same in every direction except possibly one, is a key property that motivates much of the paper. In Section 3 we describe the estimation procedure of the parameters associated to the directional regularity approach. Section 4 studies the theoretical properties of our Algorithm, where non-asymptotic bounds are provided. Section 5 describes an extensive simulation study that illustrates the good finite sample properties of our Algorithms. Section 6 discusses additional properties of our Algorithm, such as computational considerations and other interesting insights. Section 7 describes some natural applications of our directional regularity approach, including surface smoothing and anisotropic detection. All the proofs are relegated to the Supplementary Material, which also contain additional simulation results.
@@
Figure 1: In general, the worst regularity $H_{1}$ will be obtained on each direction of canonical basis, leading to isotropy (here $H_{1}<H_{2})$. This can lead to slower rates of convergence. A change-of-basis from $(\mathbf{e_{1}},\mathbf{e_{2}})$ to $(\mathbf{u_{1}},\mathbf{u_{2}})$ enables one to obtain an anisotropic process. Locating the basis ($\mathbf{u_{1}},\mathbf{u_{2}}$) is equivalent to locating the angle between $\mathbf{u_{1}}$ and $\mathbf{e_{1}}$.
@@ 2.1 Data setting and definition
where $\varepsilon^{(j)}$ are independent, centered errors with variance $\sigma^{2}$. We suppose without loss of generality that the design points $\mathbf{t}_{m}$ are observed on the canonical basis, and that $\mathcal{T}=(0,1)^{2}$. The random design case with heteroscedastic errors, together with other possible extensions, are discussed in Section 6.
@@
where $G_{\mathbf{u}}(\mathbf{t},\Delta)\stackrel{{\scriptstyle\Delta\rightarrow 0}}{{=}}o\left(\Delta^{2H_{\mathbf{u}}(\mathbf{t})}\right)$ for each fixed $\mathbf{t}$. The map
@@ 2.2 Examples
Let $H\in(0,1)$ be the Hurst parameter of the fractional Brownian motion (fBm) $\{B^{H}(t),t\in(0,1)\}$. $B^{H}$ is a centered Gaussian process with covariance function
@@
Example 2 (Sum and product of two independent Ornstein-Uhlenbeck processes).
@@ 3.1 Estimating equations
Let $(\mathbf{e_{1}},\mathbf{e_{2}})$ be the canonical basis of $\mathcal{T}$, and $(\mathbf{u_{1}},\mathbf{u_{2}})$ be orthonormal basis vectors such that $|H_{\mathbf{u}_{1}}(\mathbf{t})-H_{\mathbf{u}_{2}}(\mathbf{t})|>0$. Let $\alpha(\mathbf{t})\in[0,\pi)$ be the solution of the problem (5). Denote $H_{1}(\mathbf{t})=H_{\mathbf{u_{1}}}(\mathbf{t})$ and $H_{2}(\mathbf{t})=H_{\mathbf{u_{2}}}(\mathbf{t})$. The estimating equation for $\alpha$ is found in Proposition 1. Denote $a\wedge b=\min\{a,b\}$, for $a,b\in\mathbb{R}$, and $a\vee b=\max\{a,b\}$. Finally, we use #$\#\mathcal{S}$ to denote the cardinality of a finite set $\mathcal{S}$.
@@ 3.3 Correction for the remainder term
Figure 2: Plot showing the remainder term for different angles $\alpha$. $\mathcal{F}(\alpha)\rightarrow\infty$ as $\alpha\rightarrow\pi/2$. Going closer to the boundary 00 and $\pi/2$ can dramatically affect the estimation of the true angles.
@@ 4 Theoretical Properties
We have
$\left|H_{\mathbf{u}(\beta)}-H_{\mathbf{u}(\beta)}(\Delta)\right|=O\left(\Delta^{\zeta}\right).$ (34)
Suppose that Assumption (H3) holds. Then for any $\Delta$ sufficiently small, and for any pair of angles $(\beta_{1},\beta_{2})\in[0,2\pi]$, the following bound holds:
$\left|H_{\mathbf{u}(\beta_{1})}(\Delta)-H_{\mathbf{u}(\beta_{2})}(\Delta)\right|\leq$ $\mathfrak{a}\left\{\frac{2^{1+\underline{H}-2H_{\mathbf{u}(\beta_{2})}}}{L_{\mathbf{u}(\beta_{2})}(\mathbf{t})}\Delta^{2(\underline{H}-H_{\mathbf{u}(\beta_{2})})}+\frac{2^{1-\underline{H}}}{L_{\mathbf{u}(\beta_{1})}(\mathbf{t})}\Delta^{2(\underline{H}-H_{\mathbf{u}(\beta_{1})})}\right\}$ (35)
$\times$ $(\beta_{1}-\beta_{2})^{2\underline{H}}\times\{1+O\left(\Delta^{\zeta}\right)\}.$ (36)
@@
we have for any $r\in[0,\underline{H})$
$\mathbb{P}\left(\sup_{\beta\in[0,2\pi)}|\widehat{H}_{\mathbf{u}(\beta)}(\Delta)-H_{\mathbf{u}(\beta)}(\Delta)|\geq 2\eta\right)\leq$ $4\exp\left(-\frac{1}{4}\times\frac{\eta^{2}N\Delta^{4\underline{H}}}{2\mathfrak{b}+\mathfrak{B}\eta\Delta^{2\underline{H}}}\right)$ (38)
$+$ $4\exp\left(-\frac{1}{8}\times\frac{\eta^{2}\Delta^{4\underline{H}}N}{8\mathfrak{a}_{1}+\mathfrak{A}_{1}\eta\Delta^{2\underline{H}}}\right),$ (39)
@@
$\mathfrak{A}_{1}=4\max\left\{\frac{\mathfrak{A}}{(2M_{0})^{\underline{H}}},4\mathfrak{G}\right\},\qquad\mathfrak{a}_{1}=\frac{\mathfrak{a}}{(2M_{0})^{2\underline{H}}}+2^{4}\sigma^{2},$
@@
$\mathfrak{B}=\left(\frac{\sqrt{2}2^{\underline{H}}\mathfrak{A}}{\log(2)}\right)^{2}(\underline{H}-r)^{-3}\left\{\sqrt{2}M_{0}^{-1/2}+\Delta\right\}^{2r},\qquad\mathfrak{b}=\mathfrak{B}^{2}\left(\frac{\mathfrak{a}}{2\mathfrak{A}^{2}}\right)^{\frac{\underline{H}-r}{1-\underline{H}+r}}.$ (40)
The condition on $\eta$ in (37) represents the bias term that arises from performing denoised interpolation in order to compute the mean-squared variations. This bias term is an intrinsic feature of the “discretely observed" setup, where the surfaces are observed only at a finite number of discrete points, instead of in continuous time.
@@
Suppose that assumptions of Proposition 3 hold true, and set
@@
Moreover, we assume that a non-negative constant $\mathfrak{e}>0$ (independent of $M_{0}$ and $N$) exist such that
@@
$\sup_{\beta\in[0,2\pi)}|\widehat{H}_{\mathbf{u}(\beta)}(\Delta)-H_{\mathbf{u}(\beta)}(\Delta)|\\ =O_{\mathbb{P}}\left(\max\left\{M_{0}^{-\frac{1}{2}\underline{H}}\exp\left((2\overline{H}-\underline{H})\log(M_{0}^{\xi})\right),N^{-\frac{1}{2}}\exp\left(2\underline{H}\log(M_{0}^{\xi})\right)\right\}\right).$ (43)
The condition (42) requires that $M_{0}$ falls within two power of $N$, equivalently, the converse hold true. The choice of $\Delta$ in (41) is done is such way that $\log(M_{0})^{-b}\gg\Delta(M_{0})\gg M_{0}^{-b}$ for any fixed $b>0$, this choice can be found also in Golovkine et al., (2023) with $\xi=1/3$. the rate in (43) converge to 00 since $\Delta$ is chosen to be larger than any negative power of $M_{0}$, and keeping (42) in mind.
@@
The condition on $\varepsilon$ in (44) represent the reminder term in Proposition 1 combined with the condition (37) on $\eta$ in Proposition 3, such condition on $\varepsilon$ can be satisfied by choosing $\Delta$ as in (41) provided that $M_{0}$ is large enough. The probability bound (45) converge to zero as long as we choose $\Delta$ as in (41) and the condition (42) is satisfied.
@@
Suppose that assumptions of Proposition 3 hold true. Moreover assume that for any $k\in\{1,\dots,K_{0}\}$ we have
$\Delta_{k}\underset{\Delta\rightarrow 0}{=}o\left(\Delta^{\frac{\zeta\wedge\{\overline{H}-\underline{H}\}}{\overline{H}-\underline{H}}\underline{H}}\right).$ (46)
Then for any $\varepsilon$ such that $\Delta^{\zeta\wedge\{\overline{H}-\underline{H}\}}\ll\varepsilon$, we have for $M_{0}$ sufficiently large
$\mathbb{P}\left(|\widehat{\alpha}-\alpha|\geq 2\varepsilon\right)\leq\mathbb{P}\left(|\widehat{g(\alpha,\Delta)}-g(\alpha)|\geq\varepsilon\right)\\ +\mathbb{P}\left(|\widehat{g(\alpha,\Delta)}-g(\alpha)|^{2\underline{H}}\geq\frac{\underline{L}(\overline{H}-\underline{H})}{4\mathfrak{a}2^{3-\underline{H}}K_{0}}\right)\\ +\mathbb{P}\left(\sup_{\beta\in[0,2\pi)}|\widehat{H}_{\mathbf{u}(\beta)}(\Delta)-H_{\mathbf{u}(\beta)}(\Delta)|\geq\frac{\overline{H}-\underline{H}}{8K_{0}}\right).$ (47)
The convergence of the RHS in (47) can be guaranteed following the discussion provided for Proposition 3, and Proposition 4 and Corollary 1. The main message of Theorem 1 is that, in order to identify the right angle that gives the maximal regularity, we need to choose the values in the grid $\mathbf{\Delta}$ introduced in (23) to have a slower rate of decrease than the $\Delta$ used to estimate the function $g$ in (19). This is because, in (46), we have $(\zeta\wedge\{\overline{H}-\underline{H}\})\{\overline{H}-\underline{H}\}^{-1}\underline{H}<1$.

### v3

@@ Abstract
We introduce directional regularity, a new definition of anisotropy for multivariate functional data. Instead of taking the conventional view which determines anisotropy as a notion of smoothness along a dimension, directional regularity additionally views anisotropy through the lens of directions. We show that faster rates of convergence can be obtained through a change-of-basis by adapting to the directional regularity of a multivariate process. An algorithm for the estimation and identification of the change-of-basis matrix is constructed, made possible due to the replication structure of functional data. Non-asymptotic bounds are provided for our algorithm, supplemented by numerical evidence from an extensive simulation study. Possible applications of the directional regularity approach are discussed, and we advocate its consideration as a standard pre-processing step in multivariate functional data analysis.
@@ 1 Introduction
Functional data analysis (FDA) is a burgeoning field that deals with modern, complex data, such as those collected by sensors. It has proved to be successful in a wide-ranging number of applications, such as medical, economics and sports data (see for example, Ramsay and Silverman, (2005), Horváth and Kokoszka, (2012), Kokoszka and Reimherr, (2017)). A distinctive feature of FDA is that the datum are functions, defined over a domain $\mathcal{T}$, such as curves, surfaces and their $d$-dimensional equivalents. Typically modelled as a stochastic process in the space of squared integrable functions, the smoothness of the underlying function plays a critical role in the convergence rates (see Tsybakov, (2009)).
Since all supposedly functional data are observed on a discrete grid of points, smoothing methods (such as splines, local-polynomial smoothers, basis expansions etc) are often a necessary pre-processing step. The optimal choice of smoothing parameters depend on the regularity of the process. In higher dimensions, anisotropy affects the effective smoothness, and can be taken advantage of to get better rates of convergence through the choice of the smoothing parameters. For example, when using kernel smoothing methods, a larger bandwidth can be chosen along the dimension with higher regularity. However, anisotropy is usually determined only using the canonical basis. In this work, we show that (see Lemma 1) this can result in misleading conclusions on the effective smoothness.
In FDA, the notion of smoothness can be more subtle that the traditional non-parametric statistics literature, due to the stochastic nature of the random function. In probability theory, a rigorous notion of smoothness can be defined, for example, through the Kolmogorov-Chentsov theorem (see Revuz and Yor, (1999, Theorem 2.1)). Despite being mathematically elegant, it is not convenient for the purposes of adaptive plug-in estimation, since the regularity is defined in the form of an inequality. A recent line of work in FDA has attempted to characterize the regularity of the random function by viewing it as a stochastic process $X$, with a characterization in the form of an approximate equality:
$\mathbb{E}[\{X(u)-X(v)\}^{2}]\approx L^{2}_{t}|u-v|^{2H_{t}},\qquad\forall u\leq t\leq v,\quad t\in\mathcal{T}\subset\mathbb{R}.$
@@
See for example Golovkine et al., (2022), Wang et al., (2024), Kassi et al., (2023), Maissoro et al., (2024). Building upon this recent body of work, we aim to extend this notion of regularity to the multivariate case, where $\mathcal{T}$ is an open, bounded subset of $\mathbb{R}^{2}$. In the process of doing so, we consider a new notion of anisotropy in FDA that takes into account its directional nature, called directional regularity (see Definition 1). Through several examples, we demonstrate that performing a change-of-basis can increase the effective smoothness of a process along the canonical basis. This can lead to faster rates of convergence for ubiquitous tasks such as smoothing in FDA. This change-of-basis matrix is intimately related to the notion of directional regularity that we introduce in this paper. An algorithm is proposed to estimate this change-of-matrix, which satisfies the following important properties: (i) allows one to possibly obtain faster rates of convergence on the canonical basis; (ii) be applicable to a wide variety of tasks, such as smoothing and anisotropic detection, amongst others; (iii) be computationally feasible.
Considering and adapting to the directional nature of anisotropy has emerged in a relatively recent body of work in statistical learning, called structural adaptation (Lepski, (2015)). See Samarov and Tsybakov, (2004), Lepski and Rebelles, (2020), Ammous et al., (2024) for some examples in the density model. In spatial data, this has been considered under the umbrella of fractal analysis, for example in the seminal work of Davies and Hall, (1999). Identifying the spatial anisotropy is a natural goal in several applications such as climatology, with an example being Swiss rainfall data (Bernardi et al., (2018)).
@@
The remainder of this paper is organized as follows. In Section 2 we give a formal definition of directional regularity, which is elucidated through Lemma 1 and several examples. Lemma 1, which states that the regularity of a surface is the same in every direction except possibly one, is a key property that motivates much of the paper. In Section 3 we describe the estimation Algorithm for the directional regularity, and the construction of the change-of-basis matrix. Section 5 explores the finite sample properties of our Algorithm using a simulation study. Section 6 discusses additional properties of our Algorithm, such as computational considerations and other interesting insights. Section 7 describes some natural applications of our directional regularity approach, including surface smoothing and anisotropic detection.
@@
Figure 1: In general, the worst regularity $H_{1}$ will be obtained on the canonical basis, leading to isotropy (here $H_{1}<H_{2})$. This can lead to slower rates of convergence. A change-of-basis from $(\mathbf{e_{1}},\mathbf{e_{2}})$ to $(\mathbf{u_{1}},\mathbf{u_{2}})$ enables one to obtain an anisotropic process. Locating the basis ($\mathbf{u_{1}},\mathbf{u_{2}}$) is equivalent to locating the angle between $\mathbf{u_{1}}$ and $\mathbf{e_{1}}$.
@@ 2.1 Data setting and definition
where $\varepsilon^{(j)}$ are independent, centered errors with variance $\sigma^{2}$. We suppose without loss of generality that the design points $\mathbf{t}_{m}$ are observed on the canonical basis, and that $\mathcal{T}=(0,1)^{2}$. The random design case is briefly discussed in Section 6.
@@
where $G_{\mathbf{u}}(\mathbf{t},\Delta)=o\left(\Delta^{2H_{\mathbf{u}}(\mathbf{t})}\right)$, as $\Delta\rightarrow 0$, for each fixed $\mathbf{t}$. The map
@@ 2.2 Examples
] Let $H\in(0,1)$ be the Hurst parameter of the fractional Brownian motion (fBm) $\{B^{H}(t),t\in(0,1)\}$. $B^{H}$ is a centered Gaussian process with covariance function
@@
Example 2 (Sums and products of two independent Ornstein-Uhlenbeck processes).
@@ 3.1 Estimating equations
Let $(\mathbf{e_{1}},\mathbf{e_{2}})$ be the canonical basis of $\mathcal{T}$, and $(\mathbf{u_{1}},\mathbf{u_{2}})$ be orthonormal basis vectors such that $|H_{\mathbf{u}_{1}}(\mathbf{t})-H_{\mathbf{u}_{2}}(\mathbf{t})|>0$. Let $\alpha(\mathbf{t})\in[0,\pi)$ be the solution of the problem (4). Denote $H_{1}(\mathbf{t})=H_{\mathbf{u_{1}}}(\mathbf{t})$ and $H_{2}(\mathbf{t})=H_{\mathbf{u_{2}}}(\mathbf{t})$. The estimating equation for $\alpha$ is found in Proposition 1. Denote $a\wedge b=\min\{a,b\}$, for $a,b\in\mathbb{R}$, and $a\vee b=\max\{a,b\}$. Let ${.}\mathbbm{1}\{.\}$ be the indicator function, and #$\#\mathcal{S}$ denote the cardinality of a finite set $\mathcal{S}$.
@@ 3.3 Correction for the remainder term
Figure 2: Plot showing the remainder term for different angles $\alpha$. $\mathcal{F}(\alpha)\rightarrow\infty$ as $\alpha\rightarrow\pi/2$, for the sum of fBms as presented in Example 1 and $\Delta=M_{0}^{-1/4}$. Going closer to the boundary 00 and $\pi/2$ can dramatically affect the estimation of the true angles.
@@ 4 Theoretical Properties
It holds that
$\left|H_{\mathbf{u}(\beta)}-H_{\mathbf{u}(\beta)}(\Delta)\right|=O\left(\Delta^{\zeta}\right).$
Suppose that Assumption (3) holds. For any pair of angles $(\beta_{1},\beta_{2})\in[0,2\pi)$, and for any $\Delta$ sufficiently small, let $L_{\mathbf{u}(\beta_{i})}=(\#\widetilde{\mathcal{T}})^{-1}\sum_{t\in\widetilde{\mathcal{T}}}L_{\mathbf{u}(\beta_{i})}(\mathbf{t}),i=1,2$. Then, it holds that
$\frac{\left|H_{\mathbf{u}(\beta_{1})}(\Delta)-H_{\mathbf{u}(\beta_{2})}(\Delta)\right|}{(\beta_{1}-\beta_{2})^{2\underline{H}}}\leq\mathfrak{a}\left\{\frac{2^{1+\underline{H}-2H_{\mathbf{u}(\beta_{2})}}}{L_{\mathbf{u}(\beta_{2})}}\Delta^{2(\underline{H}-H_{\mathbf{u}(\beta_{2})})}+\frac{2^{1-\underline{H}}}{L_{\mathbf{u}(\beta_{1})}}\Delta^{2(\underline{H}-H_{\mathbf{u}(\beta_{1})})}\right\}\left\{1+O\left(\Delta^{\zeta}\right)\right\}.$ (33)
@@
we have for any $r\in[0,\underline{H})$,
$\mathbb{P}\left(\sup_{\beta\in[0,2\pi)}|\widehat{H}_{\mathbf{u}(\beta)}(\Delta)-H_{\mathbf{u}(\beta)}(\Delta)|\geq 2\eta\right)\leq$ $4\exp\left(\frac{-\eta^{2}N\Delta^{4\underline{H}}}{8\mathfrak{b}+4\mathfrak{B}\eta\Delta^{2\underline{H}}}\right)+4\exp\left(\frac{-\eta^{2}\Delta^{4\underline{H}}N}{64\mathfrak{a}_{1}+8\mathfrak{A}_{1}\eta\Delta^{2\underline{H}}}\right),$
@@
$\mathfrak{A}_{1}=\frac{4\mathfrak{A}}{(2M_{0})^{\underline{H}}}\vee 16\mathfrak{G},\qquad\mathfrak{a}_{1}=\frac{\mathfrak{a}}{(2M_{0})^{2\underline{H}}}+2^{4}\sigma^{2},\qquad\mathfrak{b}=\mathfrak{B}^{2}\left(\frac{\mathfrak{a}}{2\mathfrak{A}^{2}}\right)^{\frac{\underline{H}-r}{1-\underline{H}+r}},$
@@
$\mathfrak{B}=\left(\frac{\sqrt{2}2^{\underline{H}}\mathfrak{A}}{\log(2)}\right)^{2}(\underline{H}-r)^{-3}\left\{\sqrt{2}M_{0}^{-1/2}+\Delta\right\}^{2r}.$
The condition on $\eta$ in (34) represents the bias term that arises from performing denoised interpolation in order to compute the mean-squared variations. This bias term is an intrinsic feature of the “discretely observed" setup, where the surfaces are observed only at a finite number of discrete points, instead of in continuous time. The following corollary is a direct consequence of Proposition 3, with an appropriate choice of $\Delta$.
@@
Suppose that the assumptions of Proposition 3 hold true, and set
@@
Moreover, we assume that a non-negative constant $\mathfrak{e}>0$ (independent of $M_{0}$ and $N$) exists such that
@@
$\sup_{\beta\in[0,2\pi)}|\widehat{H}_{\mathbf{u}(\beta)}(\Delta)-H_{\mathbf{u}(\beta)}(\Delta)|=O_{\mathbb{P}}\left(\frac{\exp\left((2\overline{H}-\underline{H})\log(M_{0})^{\xi}\right)}{\sqrt{M_{0}}^{\underline{H}}}\vee\frac{\exp\left(2\underline{H}\log(M_{0})^{\xi}\right)}{\sqrt{N}}\right).$ (37)
The condition (36) requires that $M_{0}$ falls within two powers of $N$, or equivalently, that the converse holds true. The choice of $\Delta$ in (35) is done in such a way that $\log(M_{0})^{-b}\gg\Delta\gg M_{0}^{-b}$ for any fixed $b>0$. This choice can also be found in Golovkine et al., (2022) with $\xi=1/3$. The rate terms in (37) converges to 00 since the choice of $\Delta$ is larger than any negative power of $M_{0}$, and keeping (36) in mind.
@@
The condition on $\varepsilon$ in (38) represents the remainder term in Proposition 1. Combined with the condition (34) on $\eta$ in Proposition 3, such a condition on $\varepsilon$ can be satisfied by choosing $\Delta$ as in (35), provided that $M_{0}$ is sufficiently large. The probability bound in (39) converges to zero as long as we choose $\Delta$ as in (35), and the condition (36) is satisfied.
@@
Suppose that the assumptions of Proposition 3 hold true. Moreover assume that for any $k\in\{1,\dots,K_{0}\}$, we have
$\Delta_{k}=o\left(\Delta^{\frac{\zeta\wedge\{\overline{H}-\underline{H}\}}{\overline{H}-\underline{H}}\underline{H}}\right),\qquad\text{as }\Delta\rightarrow 0.$ (40)
Then for any $\varepsilon$ such that $\Delta^{\zeta\wedge\{\overline{H}-\underline{H}\}}\ll\varepsilon$, we have for $M_{0}$ sufficiently large,
$\mathbb{P}\left(|\widehat{\alpha}-\alpha|\geq 2\varepsilon\right)\leq\mathbb{P}\left(\sup_{\beta\in[0,2\pi)}|\widehat{H}_{\mathbf{u}(\beta)}(\Delta)-H_{\mathbf{u}(\beta)}(\Delta)|\geq\frac{\overline{H}-\underline{H}}{8K_{0}}\right)\\ +\mathbb{P}\left(|\widehat{g(\alpha,\Delta)}-g(\alpha)|\geq\varepsilon\right)+\mathbb{P}\left(|\widehat{g(\alpha,\Delta)}-g(\alpha)|^{2\underline{H}}\geq\frac{\underline{L}(\overline{H}-\underline{H})}{4\mathfrak{a}2^{3-\underline{H}}K_{0}}\right).$ (41)
The convergence of the RHS in (41) can be guaranteed following the discussion of Propositions 3, 4 and Corollary 1. The main message of Theorem 1 is that in order to identify the right angle which gives the maximal regularity, we need to choose the values in the grid $\boldsymbol{\Delta}$, introduced in (22), to have a slower rate of decrease than the $\Delta$ used to estimate the function $g$ in (18). This is because in (40), we have $(\zeta\wedge\{\overline{H}-\underline{H}\})\{\overline{H}-\underline{H}\}^{-1}\underline{H}<1$.